# Debiased Inference for Dynamic Nonlinear Panels with Multi-dimensional Heterogeneities

### v4

@@ A Proofs
$\sqrt{NT}\triangledown_{\theta}L(\widehat{\theta}_{R})+[J(\widehat{\theta}_{R})]^{\prime}\sqrt{NT}\widehat{\lambda}$ $=0,$ (A.27)
$\sqrt{NT}R(\widehat{\theta}_{R})$ $=0.$ (A.28)
Following a similar proof as in Corollary 1, we have $\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})=O_{\mathbb{P}}\left(1\right)$. Furthermore, $\sqrt{NT}\widehat{\lambda}=O_{\mathbb{P}}(1)$ as implied by (A.27) since $\sqrt{NT}\triangledown_{\theta}L(\theta_{0})=O_{\mathbb{P}}(1)$ and $J(\theta_{0})$ is of full row rank. By Theorem 2, a mean value expansion of $\triangledown_{\theta}L(\widehat{\theta}_{R})$ around $\theta_{0}$ is
@@
where each component of $\overline{\theta}$ lies between the corresponding components of $\widehat{\theta}_{R}$ and $\theta_{0}.$ Combining this with (A.27) yields that
$\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})+[J(\theta_{0})]^{\prime}\sqrt{NT}\widehat{\lambda}=-\sqrt{NT}\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}(1).$ (A.29)
Under $H_{0}$ and from (A.28), we also obtain
00 0 $=\sqrt{NT}R(\widehat{\theta}_{R})$
@@
$=J(\theta_{0})\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})+o_{\mathbb{P}}(1).$ (A.30)
Now, combining (A.29) and (A.30), we have
@@
$\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})=-\left(\{\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\}^{-1}-\Xi\left(\theta_{0}\right)\right)\sqrt{NT}\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}\left(1\right).$ (A.31)
The conclusion follows by subtracting (A.25) in Corollary 1 from both sides of (A.31).
Finally, we proof the asymptotic distribution of $\widehat{\xi}_{LR},\widehat{\xi}_{LM},\widehat{\xi}_{Wald}$ below.
Proof of Corollary 2. Let $\widehat{\xi}:=-NT\triangledown_{\theta^{\prime}}l(\theta_{0})\Xi\left(\theta_{0}\right)\triangledown_{\theta}l(\theta_{0})$ for $\Xi\left(\theta_{0}\right)$ defined in Lemma A.4. We first show that
$\widehat{\xi}_{LR}=\widehat{\xi}+o_{\mathbb{P}}(1),\qquad\widehat{\xi}_{LM}=\widehat{\xi}+o_{\mathbb{P}}(1),\qquad\widehat{\xi}_{Wald}=\widehat{\xi}+o_{\mathbb{P}}(1).$ (A.32)
@@
Since $\triangledown_{\theta\theta^{\prime}}L(\dot{\theta})-\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})]=o_{\mathbb{P}}(1)$, by Lemma A.4, we further conclude that
@@
[…]
@@
B Extra Simulation Results
B.1 Logit Model
Table B.1: Empirical Size of LM and Wald Tests for Logit Model, Design 1
(a) LM Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $15\%$ $9\%$ $18\%$ $10\%$
$(60,60)$ $15\%$ $5\%$ $17\%$ $7\%$
$(90,90)$ $18\%$ $6\%$ $16\%$ $5\%$
(b) Wald Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $15\%$ $9\%$ $18\%$ $10\%$
$(60,60)$ $15\%$ $5\%$ $17\%$ $7\%$
$(90,90)$ $18\%$ $6\%$ $16\%$ $5\%$
[…]
Notes: Logit model with $\varepsilon_{it}$ standard logistic. Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$. The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level $5\%$. The number of replications is $1000$.
Table B.2: Empirical Power of LM and Wald Tests for Logit Model using $L(\theta)$, Design 1
(a) LM Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $44\%$ $32\%$ $22\%$ $10\%$ $14\%$ $21\%$
$(60,60)$ $86\%$ $62\%$ $37\%$ $25\%$ $50\%$ $74\%$
$(90,90)$ $100\%$ $93\%$ $65\%$ $50\%$ $84\%$ $98\%$
Dynamic Model
$(30,30)$ $48\%$ $32\%$ $20\%$ $18\%$ $28\%$ $43\%$
$(60,60)$ $94\%$ $73\%$ $40\%$ $45\%$ $75\%$ $94\%$
$(90,90)$ $100\%$ $98\%$ $76\%$ $78\%$ $98\%$ $100\%$
(b) Wald Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $44\%$ $32\%$ $22\%$ $10\%$ $14\%$ $21\%$
$(60,60)$ $86\%$ $62\%$ $37\%$ $25\%$ $50\%$ $74\%$
$(90,90)$ $100\%$ $93\%$ $65\%$ $50\%$ $84\%$ $98\%$
Dynamic Model
$(30,30)$ $47\%$ $31\%$ $20\%$ $18\%$ $28\%$ $43\%$
$(60,60)$ $94\%$ $73\%$ $39\%$ $45\%$ $75\%$ $94\%$
$(90,90)$ $100\%$ $98\%$ $76\%$ $78\%$ $98\%$ $100\%$
[…]
Notes: Logit model with $\varepsilon_{it}$ standard logistic. Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$. The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$. The number of replications is $1000$.
Table B.3: Comparisons of Uncorrected and Corrected Logit Estimates, Design 2
@@ B.1 Logit Model
Static Model
$\widehat{\beta}$ $0.454$ $1.201$ $0.100$ $0.132$ $0.060$ $0.071$
$\widehat{\beta}_{L}$ $0.101$ $0.199$ $0.020$ $0.056$ $0.010$ $0.035$
Dynamic Model
$\widehat{\beta}$ $0.527$ $1.583$ $0.109$ $0.212$ $0.059$ $0.071$
$\widehat{\beta}_{L}^{(1)}$ $0.117$ $0.204$ $0.024$ $0.058$ $0.009$ $0.038$
$\widehat{\beta}_{L}^{(2)}$ $0.126$ $0.278$ $0.025$ $0.059$ $0.009$ $0.041$
$\widehat{\rho}$ $-0.110$ $0.241$ $-0.055$ $0.110$ $-0.037$ $0.071$
$\widehat{\rho}_{L}^{(1)}$ $-0.041$ $0.192$ $-0.017$ $0.090$ $-0.010$ $0.059$
$\widehat{\rho}_{L}^{(2)}$ $-0.035$ $0.196$ $-0.012$ $0.091$ $-0.007$ $0.059$
@@
Notes: Logit model with $\varepsilon_{it}$ standard logistic. Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$. For the static model, $\widehat{\beta}$ and $\widehat{\beta}_{L}$ are, respectively, the uncorrected and corrected estimates of $\beta_{0}=0.5$. For the dynamic model $(\widehat{\beta},\widehat{\rho})^{\prime}$ is the uncorrected estimate of $(\rho_{0},\beta_{0})^{\prime}=(0.5,0.5)^{\prime}$ from and $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ is the corrected estimate with truncation parameter $\tau\in(1,2)$. The number of replications is $1000$.
Table B.4: Empirical Size of LR, LM, and Wald Tests for Logit Model, Design 2
(a) LR Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $47\%$ $20\%$ $43\%$ $18\%$
$(60,60)$ $43\%$ $10\%$ $38\%$ $9\%$
$(90,90)$ $42\%$ $9\%$ $34\%$ $7\%$
(b) LM Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $42\%$ $18\%$ $38\%$ $16\%$
$(60,60)$ $42\%$ $9\%$ $37\%$ $9\%$
$(90,90)$ $41\%$ $9\%$ $34\%$ $7\%$
(c) Wald Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $32\%$ $15\%$ $27\%$ $15\%$
$(60,60)$ $41\%$ $9\%$ $36\%$ $8\%$
$(90,90)$ $42\%$ $9\%$ $34\%$ $7\%$
@@
Notes: Logit model with $\varepsilon_{it}$ standard logistic. Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$. The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level $5\%$. The number of replications is $1000$.
Table B.5: Empirical Power of LR, LM, and Wald Tests for Logit Model using $L(\theta)$, Design 2
(a) LR Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $79\%$ $62\%$ $45\%$ $12\%$ $16\%$ $26\%$
$(60,60)$ $100\%$ $95\%$ $73\%$ $40\%$ $75\%$ $93\%$
$(90,90)$ $100\%$ $100\%$ $94\%$ $81\%$ $98\%$ $100\%$
Dynamic Model
$(30,30)$ $75\%$ $58\%$ $41\%$ $19\%$ $26\%$ $35\%$
$(60,60)$ $100\%$ $94\%$ $69\%$ $44\%$ $80\%$ $96\%$
$(90,90)$ $100\%$ $100\%$ $93\%$ $86\%$ $99\%$ $100\%$
(b) LM Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $77\%$ $60\%$ $42\%$ $12\%$ $17\%$ $28\%$
$(60,60)$ $99\%$ $95\%$ $72\%$ $41\%$ $77\%$ $93\%$
$(90,90)$ $100\%$ $100\%$ $94\%$ $82\%$ $98\%$ $100\%$
Dynamic Model
$(30,30)$ $73\%$ $56\%$ $38\%$ $19\%$ $27\%$ $36\%$
$(60,60)$ $100\%$ $93\%$ $68\%$ $45\%$ $80\%$ $94\%$
$(90,90)$ $100\%$ $100\%$ $93\%$ $86\%$ $99\%$ $99\%$
(c) Wald Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $75\%$ $57\%$ $39\%$ $10\%$ $17\%$ $29\%$
$(60,60)$ $100\%$ $95\%$ $72\%$ $41\%$ $77\%$ $95\%$
$(90,90)$ $100\%$ $100\%$ $94\%$ $81\%$ $98\%$ $100\%$
Dynamic Model
$(30,30)$ $69\%$ $50\%$ $34\%$ $18\%$ $26\%$ $36\%$
$(60,60)$ $100\%$ $93\%$ $68\%$ $45\%$ $81\%$ $97\%$
$(90,90)$ $100\%$ $100\%$ $92\%$ $86\%$ $99\%$ $100\%$
@@
Notes: Logit model with $\varepsilon_{it}$ standard logistic. Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$. The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$. The number of replications is $1000$.
B.2 Probit Model
@@
Table B.6: Comparisons of Uncorrected and Corrected Probit Estimates, Design 1

### v5

@@ E Proofs
$\sqrt{NT}\triangledown_{\theta}L(\widehat{\theta}_{R})+[J(\widehat{\theta}_{R})]^{\prime}\sqrt{NT}\widehat{\lambda}$ $=0,$ (E.27)
$\sqrt{NT}R(\widehat{\theta}_{R})$ $=0.$ (E.28)
Following a similar proof as in Corollary 1, we have $\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})=O_{\mathbb{P}}\left(1\right)$. Furthermore, $\sqrt{NT}\widehat{\lambda}=O_{\mathbb{P}}(1)$ as implied by (E.27) since $\sqrt{NT}\triangledown_{\theta}L(\theta_{0})=O_{\mathbb{P}}(1)$ and $J(\theta_{0})$ is of full row rank. By Theorem 2, a mean value expansion of $\triangledown_{\theta}L(\widehat{\theta}_{R})$ around $\theta_{0}$ is
@@
where each component of $\overline{\theta}$ lies between the corresponding components of $\widehat{\theta}_{R}$ and $\theta_{0}.$ Combining this with (E.27) yields that
$\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})+[J(\theta_{0})]^{\prime}\sqrt{NT}\widehat{\lambda}=-\sqrt{NT}\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}(1).$ (E.29)
Under $H_{0}$ and from (E.28), we also obtain
$0$ $=\sqrt{NT}R(\widehat{\theta}_{R})$
@@
$=J(\theta_{0})\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})+o_{\mathbb{P}}(1).$ (E.30)
Now, combining (E.29) and (E.30), we have
@@
$\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})=-\left(\{\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\}^{-1}-\Xi\left(\theta_{0}\right)\right)\sqrt{NT}\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}\left(1\right).$ (E.31)
The conclusion follows by subtracting (E.25) in Corollary 1 from both sides of (E.31).
Below, we proof the asymptotic distribution of $\widehat{\xi}_{LR},\widehat{\xi}_{LM},\widehat{\xi}_{Wald}$ below.
Proof of Corollary 2. Let $\widehat{\xi}:=-NT\triangledown_{\theta^{\prime}}l(\theta_{0})\Xi\left(\theta_{0}\right)\triangledown_{\theta}l(\theta_{0})$ for $\Xi\left(\theta_{0}\right)$ defined in Lemma E.4. We first show that
$\widehat{\xi}_{LR}=\widehat{\xi}+o_{\mathbb{P}}(1),\qquad\widehat{\xi}_{LM}=\widehat{\xi}+o_{\mathbb{P}}(1),\qquad\widehat{\xi}_{Wald}=\widehat{\xi}+o_{\mathbb{P}}(1).$ (E.32)
@@
Since $\triangledown_{\theta\theta^{\prime}}L(\dot{\theta})-\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})]=o_{\mathbb{P}}(1)$, by Lemma E.4, we further conclude that
@@
where each component of $\ddot{\theta}$ lies between the corresponding components of $\widehat{\theta}_{R}$ and $\widehat{\theta}_{L}$. Thus,
@@
F Additional Simulation Results
F.1 Additional Results under Design 1
Table F.1: Comparisons of Uncorrected and Corrected Probit Estimates, Design 1
@@ F.1 Additional Results under Design 1
Static Probit Model
$\widehat{\beta}$ $0.175$ $0.208$ $0.069$ $0.076$ $0.044$ $0.048$
$\widehat{\beta}_{L}$ $0.067$ $0.101$ $0.016$ $0.034$ $0.008$ $0.019$
$\widehat{\beta}_{J}$ 1.966-1.966 $3.942$ 0.040-0.040 $0.101$ 0.010-0.010 $0.022$
$\widehat{\beta}_{B}$ 0.114-0.114 $0.153$ 0.013-0.013 $0.030$ 0.005-0.005 $0.018$
Dynamic Probit Model
$\widehat{\beta}$ $0.209$ $0.335$ $0.072$ $0.080$ $0.045$ $0.049$
$\widehat{\beta}_{L}^{(1)}$ $0.079$ $0.116$ $0.019$ $0.036$ $0.008$ $0.021$
$\widehat{\beta}_{L}^{(2)}$ $0.082$ $0.118$ $0.020$ $0.037$ $0.009$ $0.021$
$\widehat{\beta}_{J}$ 2.161-2.161 $3.956$ 0.077-0.077 $0.218$ 0.013-0.013 $0.025$
$\widehat{\beta}_{B}$ 0.155-0.155 $0.482$ 0.015-0.015 $0.032$ 0.006-0.006 $0.019$
$\widehat{\rho}$ 0.044-0.044 $0.126$ 0.021-0.021 $0.059$ 0.015-0.015 $0.036$
$\widehat{\rho}_{L}^{(1)}$ 0.024-0.024 $0.109$ 0.009-0.009 $0.053$ 0.007-0.007 $0.032$
$\widehat{\rho}_{L}^{(2)}$ 0.016-0.016 $0.109$ 0.004-0.004 $0.052$ 0.003-0.003 $0.032$
$\widehat{\rho}_{J}$ 0.008-0.008 $0.116$ $0.001$ $0.055$ $0.000$ $0.032$
$\widehat{\rho}_{B}$ 0.004-0.004 $0.100$ $0.000$ $0.052$ 0.001-0.001 $0.031$
@@
Table F.2: Empirical Size and Power from LR Test for Probit Model, Design 1
(a) Size
Size
$(N,T)$ Uncorrected Analytical Bootstrap
Static Probit Model
$(30,30)$ $66$ $26$ 22
$(60,60)$ $64$ $13$ 22
$(90,90)$ $66$ 88 33
Dynamic Probit Model
$(30,30)$ $61$ $25$ 22
$(60,60)$ $60$ $12$ 33
$(90,90)$ $57$ 99 22
(b) Power
Power at $\delta$ (Analytical) Power at $\delta$ (Bootstrap)
$(N,T)$ 0.2-0.2 0.1-0.1 $0.1$ $0.2$ 0.2-0.2 0.1-0.1 $0.1$ $0.2$
Static Probit Model
$(30,30)$ $99$ $80$ $16$ $58$ $85$ $24$ 0 0
$(60,60)$ $100$ $99$ $83$ $100$ $100$ $96$ 0 $56$
$(90,90)$ $100$ $100$ $100$ $100$ $100$ $100$ $15$ $100$
Dynamic Probit Model
$(30,30)$ $99$ $74$ $28$ $73$ $78$ $18$ 11 22
$(60,60)$ $100$ $99$ $88$ $100$ $100$ $94$ 44 $86$
$(90,90)$ $100$ $100$ $100$ $100$ $100$ $100$ $50$ $100$
@@
Notes: See also Table F.5 where we show that the sizes for the probit model with $\left(N,T\right)=\left(120,120\right)$ and $\left(150,150\right)$ are close to $5\%$.
Table F.3: Empirical Size and Power from LM Test for Logit and Probit Model, Design 1
Size Power at $\delta$
$(N,T)$ Uncorrected Analytical 0.2-0.2 0.1-0.1 $0.1$ $0.2$
Static Logit Model
$(30,30)$ $37$ $13$ $87$ $47$ $13$ $46$
$(60,60)$ $40$ 88 $100$ $85$ $59$ $99$
$(90,90)$ $38$ 55 $100$ $98$ $95$ $100$
Static Probit Model
$(30,30)$ $64$ $25$ $99$ $79$ $17$ $60$
$(60,60)$ $63$ $13$ $100$ $99$ $84$ $100$
$(90,90)$ $66$ 88 $100$ $100$ $100$ $100$
Dynamic Logit Model
$(30,30)$ $35$ $13$ $83$ $41$ 2020 $56$
$(60,60)$ $35$ 99 $100$ $79$ $64$ $99$
$(90,90)$ $33$ 66 $100$ $98$ $95$ $100$
Dynamic Probit Model
$(30,30)$ $58$ $22$ $99$ $73$ $28$ $75$
$(60,60)$ $58$ $12$ $100$ $99$ $89$ $99$
$(90,90)$ $56$ 99 $100$ $100$ $100$ $99$
@@
Notes: See also Table F.5 where we show that the sizes for the probit model with $\left(N,T\right)=\left(120,120\right)$ and $\left(150,150\right)$ are close to $5\%$.
Table F.4: Empirical Size and Power from Wald Test for Logit and Probit Model, Design 1
Size Power at $\delta$
$(N,T)$ Uncorrected Analytical 0.2-0.2 0.1-0.1 $0.1$ $0.2$
Static Logit Model
$(30,30)$ $37$ $13$ $87$ $46$ $12$ $45$
$(60,60)$ $40$ 88 $100$ $85$ $59$ $99$
$(90,90)$ $38$ 55 $100$ $98$ $95$ $100$
Static Probit Model
$(30,30)$ $63$ $23$ $98$ $78$ $17$ $59$
$(60,60)$ $63$ $13$ $100$ $99$ $84$ $100$
$(90,90)$ $66$ 88 $100$ $100$ $100$ $100$
Dynamic Logit Model
$(30,30)$ $34$ $12$ $82$ $40$ 2020 $56$
$(60,60)$ $35$ 99 $100$ $79$ $64$ $99$
$(90,90)$ $32$ 66 $100$ $98$ $95$ $100$
Dynamic Probit Model
$(30,30)$ $55$ 2121 $98$ $72$ $29$ $75$
$(60,60)$ $58$ $12$ $99$ $98$ $89$ $99$
$(90,90)$ $56$ 99 $100$ $100$ $100$ $100$
[…]
Notes: See also Table F.5 where we show that the sizes for the probit model with $\left(N,T\right)=\left(120,120\right)$ and $\left(150,150\right)$ are close to $5\%$.
Table F.5: Empirical Size from LR, LM, and Wald Test for Probit Model, Design 1 with $N,T=(120,120)$ and $(150,150)$
LR Test LM Test Wald Test
$(N,T)$ Uncorrected Analytical Uncorrected Analytical Uncorrected Analytical
Static Probit Model
$(120,120)$ $63$ 66 $63$ 66 $63$ 66
$(150,150)$ $64$ 55 $63$ 55 $63$ 55
Dynamic Probit Model
$(120,120)$ $59$ 88 $59$ 88 $59$ 88
$(150,150)$ $56$ 77 $56$ 77 $56$ 77
[…]
@@
Table F.6: Comparisons of Uncorrected and Corrected Logit and Probit Estimates, Design 1 with $N=90$ and $T=10$
Probit Logit Probit Logit
Bias RMSE Bias RMSE Bias RMSE Bias RMSE
Static Model, $\theta_{0}=\beta_{0}=0$ Static Model, $\theta_{0}=\beta_{0}=0.5$
$\widehat{\beta}$ $0.056$ $2.821$ $0.074$ $1.029$ $4.048$ $6.310$ $1.742$ $2.523$
$\widehat{\beta}_{L}$ $0.039$ $1.048$ $0.084$ $1.030$ $0.623$ $1.577$ $0.299$ $1.029$
Dynamic Model, $\theta_{0}=\left(0,0\right)$ Dynamic Model, $\theta_{0}=\left(0.5,0.5\right)$
$\widehat{\beta}$ $0.104$ $2.501$ $0.081$ $1.342$ $3.931$ $6.008$ $1.953$ $2.982$
$\widehat{\beta}_{L}$ $0.045$ $0.733$ $0.039$ $0.467$ $0.520$ $1.324$ $0.326$ $1.348$
$\widehat{\rho}$ 0.331-0.331 $0.354$ 0.537-0.537 $0.572$ 0.266-0.266 $0.298$ 0.465-0.465 $0.507$
$\widehat{\rho}_{L}$ 0.161-0.161 $0.197$ 0.230-0.230 $0.284$ 0.161-0.161 $0.222$ 0.240-0.240 $0.305$
[…]
Notes: $\left(N,T\right)=\left(90,10\right)$. For the dynamic model we set $\tau=1$, producing corrected estimators $\widehat{\rho}_{L}$ and $\widehat{\beta}_{L}$ for $\rho_{0}$ and $\beta_{0}$ respectively.
F.2 Heterogeneous Autoregressive Parameters
In this section, we present simulation studies for models with two-way heterogeneous autoregressive coefficient. We consider three models here: a Poission model, a logit model, and a probit model. For the Poisson, we consider a model specified according to the data-generating process $Y_{it}\sim\operatorname*{Poisson}\left(\lambda_{it}\right)$ for
$\lambda_{it}=\exp\left\{\left(\rho_{0}+\alpha_{1,i}^{0}+\gamma_{1,t}^{0}\right)Y_{it-1}+(\beta_{0}+\alpha_{2,i}^{0}+\gamma_{2,t}^{0})Z_{it}+\alpha_{3,i}^{0}+\gamma_{3,t}^{0}\right\},$
with $Y_{i0}\sim\operatorname*{Poisson}\left(\exp\left\{(\beta_{0}+\alpha_{1,i}^{0})Z_{i0}+\alpha_{2,i}^{0}\right\}\right)$, where the parameter of interests is $\theta_{0}=(\rho_{0},\beta_{0})^{\prime}=(-0.5,0.5)^{\prime}$. We set $\rho_{0}=-0.5$ to keep $Y_{it}$ from exploding as $t$ progresses. Similar to Section 4, we generate $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.04)$ for $k=1,\ldots,3$ and demean them afterwards [^10] As in Section 4, the individual effects are not normalized when the model is being estimated.. The rest of the experiment is the same as Design 1 in Section 4.
Table F.7 presents the simulation results. We see that the bias of the MLE $\widehat{\beta}$ of a static Poisson model is relative small, compared to the probit and logit models above. The biases in $\widehat{\beta}_{L}$, $\widehat{\beta}_{J}$, and $\widehat{\beta}_{B}$ are all very small, with $\widehat{\beta}_{L}$ and $\widehat{\beta}_{B}$ very similar and outperforming $\widehat{\beta}_{J}$. For the dynamic Poisson model, however, the MLE $\widehat{\rho}$ is severely biased, showing a bias of 0.204-0.204 for $\left(N,T\right)=\left(30,30\right)$. On the other hand, our bias correction procedure reduces the bias considerably, delivering $\widehat{\rho}_{L}^{(1)}$ which contains a bias of 0.077-0.077, roughly $1/3$ of the bias of the MLE. Our procedure also outperforms the jackknife and the bootstrap on $\rho$ for small samples. As the sample size increases, the biases of all estimators decrease, as expected, with the performance of our procedure and that of the bootstrap similar, outperforming the jackknife. For $\left(N,T\right)=\left(30,30\right)$, we notice that the bias of $\widehat{\rho}_{B}$ is $1.323$, which is very severe. We have double checked and this seems to be a valid result. We think this might be a coincidence and would like to emphasize that, as the sample size increases, the bootstrap is an effective bias correction device. For the hypothesis tests, we only report the results from the LR test for the dynamic model in Table F.8. For the static model, the sizes of the uncorrected, corrected, and the bootstrap LR test are all very close to the nominal level of $5\%$. The patterns of the LR test are similar to those from the probit and logit models in Section 4. For the LM and Wald test, the results are comparable to the LR test, and are, therefore, omitted here.
Table F.7: Comparisons of Uncorrected and Corrected Poisson Estimates